OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is -6, the last term is -24,576, and the common ratio is 4?

OpenStudy (goformit100):

ggggggggggggggg

OpenStudy (anonymous):

g what

OpenStudy (anonymous):

not helpful -___-

OpenStudy (anonymous):

well its easy,im 100 percent sure its math

OpenStudy (anonymous):

yeahhh & i need help finding the answer :(

OpenStudy (anonymous):

hate math lol

OpenStudy (anonymous):

gee thanks ...

OpenStudy (anonymous):

lol=)

OpenStudy (anonymous):

Common ratio of 4 means you multiply each term by 4 to get the next term.

OpenStudy (amistre64):

formula for sum of a geo seq is\[\frac{g_1(1-r^n)}{1-r}\]

OpenStudy (anonymous):

use the formula for a geometric series. Sum = (first term) * (1 - r^n)/(1 - r). Just plug in the values of the first term, r and n.

OpenStudy (anonymous):

@amistre64 can you please use numbers? i'm dyslexic & the letters are con fusing me :(

OpenStudy (anonymous):

I am not so great at this either was just trying to give you some hints cuz thought maybe that would help. You can also try looking this question up on google an they have example of one similar to this, maybe that would better help u understan... sorry chic:(

OpenStudy (anonymous):

its okay :(

OpenStudy (amistre64):

show me what you think it says ... at least that way i got something i can work with. Im trying not to just blurt out an answer

OpenStudy (anonymous):

g1 = -6 . 1-r^n = -16383

OpenStudy (anonymous):

1-r = -2

OpenStudy (anonymous):

49149 ?

OpenStudy (amistre64):

the first part is excellent; the 1-r tho ... 1 - (4) = -3 this gives us: -6(-16383)/(-3)

OpenStudy (anonymous):

oh oops.

OpenStudy (amistre64):

google calculator gives me: -32766

OpenStudy (anonymous):

thats what i got on wolfram

OpenStudy (anonymous):

i have one more. its: What is the sum of a 22-term arithmetic sequence where the first term is 57 and the last term is -27?

OpenStudy (anonymous):

the answer i got is 300. is that right?

OpenStudy (amistre64):

(first + last) times half the n

OpenStudy (amistre64):

330 is what i think at first glance

OpenStudy (amistre64):

57-27 = 30 11*30 = 330

OpenStudy (anonymous):

oh i see what i did wrong